OpenStudy (anonymous):

Mary takes 3 hours to mow the yard. It takes Sally 5 hours. How long would it take with them mowing it together?

OpenStudy (shane_b):

Sorry...phone. The easiest way to do these types of problems is to know: rate1 + rate2 = combined rate For this problem: Mary's rate = 1/3 jobs/hour Sallie's rate = 1/5 jobs/hour\[\frac{1}{3}+\frac{1}{5}=\frac{1}{x}\]Just solve for x...what do you get?

OpenStudy (anonymous):

i solved it thanx (:

OpenStudy (shane_b):

What did you end up with?

OpenStudy (anonymous):

1.875 hours

OpenStudy (shane_b):

Great :)